Hoi Chong Transport, Ltd., operates a fleet of delivery trucks in Singapore. The company has determined that if a truck is driven 162,000 kilometers during a year, the average operating cost is 13.2 cents per kilometer. If a truck is driven only 108,000 kilometers during a year, the average operating cost increases to 16.6 cents per kilometer. Required: 1. Using the high-low method, estimate the variable operating cost per kilometer and the annual fixed operating cost associated with the fleet of trucks. 2. Express the variable and fixed costs in the form Y = a + bX. 3. If a truck were driven 135,000 kilometers during a year, what total operating cost would you expect to be incurred?
Answer #1

Ans:

1) Variable cost using high low method = Difference in costs/Difference between the distance travelled
Annual cost of operating
For 162000 kms@13.2 cents = 2138400 cents
For 108000 kms@16.6 cents = 1792800 cents
Variable cost = 2138400-1792800/162000-108000/
= 345600/54000
= 6.4 cents per km or $0.064
So Fixed cost = 162000(13.2-6.4) = 162000(6.8 cents)= 1101600
Variable cost = 6.4 cents per km or $0.064
Fixed cost = 1101600

b) So, to express it in the formula where
a= fixed cost
b= Variable cost per unit
X= Number of kms
Y= $11016+0.064X

c) If 135000 kms were driven, the total cost will be
= $11016+ 0.064*135000
= $11016+ $7680
= $19656